# MVAD: A Multiple Visual Artifact Detector for Video Streaming

### v1

@@ Abstract
Visual artifacts are often introduced into streamed video content, due to prevailing conditions during content production and/or delivery. Since these can degrade the quality of the user’s experience, it is important to automatically and accurately detect them in order to enable effective quality measurement and enhancement. Existing detection methods often focus on a single type of artifact and/or determine the presence of an artifact through thresholding objective quality indices. Such approaches have been reported to offer inconsistent prediction performance and are also impractical for real-world applications where multiple artifacts co-exist and interact. In this paper, we propose a Multiple Visual Artifact Detector, MVAD, for video streaming which, for the first time, is able to detect multiple artifacts using a single framework that is not reliant on video quality assessment models. Our approach employs a new Artifact-aware Dynamic Feature Extractor (ADFE) to obtain artifact-relevant spatial features within each frame for multiple artifact types. The extracted features are further processed by a Recurrent Memory Vision Transformer (RMViT) module, which captures both short-term and long-term temporal information within the input video. The proposed network architecture is optimized in an end-to-end manner based on a new, large and diverse training database that is generated by simulating the video streaming pipeline and based on Adversarial Data Augmentation. This model has been evaluated on two video artifact databases, Maxwell and BVI-Artifact, and achieves consistent and improved prediction results for ten target visual artifacts when compared to seven existing single and multiple artifact detectors. The source code and training database will be available at https://chenfeng-bristol.github.io/MVAD/.
@@ 1 Introduction
With the significant growth in subscribers on popular streaming platforms [38], such as Netflix and Amazon Prime Video, video streaming applications are now one of the largest consumers of global Internet bandwidth. For video streaming service providers, it is essential to monitor the quality of a user’s experience during viewing. However, this process is complex because streamed content can be impacted by multiple visual artifacts introduced at different production and delivery stages including acquisition, post-production, compression, and transmission [6]. For example, source artifacts such as motion blur, darkness and graininess can be produced during acquisition; banding and aliasing may be introduced in post-production; video compression can introduce artifacts such as spatial blur and blockiness; and finally packet loss in transmission leads to transmission errors, frame dropping, or black frames. These artifacts can affect the perceived quality of streamed video content and thus degrade the quality of the user’s experience. Therefore, it is important to identify the presence of these artifacts in order to enable appropriate enhancement processes and provide system feedback.
Although video quality assessment is a well-established research area, with numerous classic [45, 47, 53, 30, 28, 29, 35] and learning-based approaches [23, 19, 21, 42, 11] developed over the past two decades, these methods can only provide a quantitative prediction of perceptual quality, rather than detect and identify the artifacts appearing in a video sequence. To achieve accurate artifact detection, some existing works focus on employing video quality assessment models and apply thresholding to determine the existence of artifacts, with MaxVQA [50] being a typical example. Other works have developed bespoke models for individual artifacts, such as CAMBI [40], BBAND [41], EFENet [57], MLDBD [58] and [48]. This latter approach is a less practical solution for real-world scenarios where multiple artifacts co-exist and interact. In a recent benchmarking experiment [10], all these methods were reported to perform poorly on a video database containing multiple source and non-source artifacts - with the average artifact detection accuracy for the best performers being only 55%.
@@
  1. 1.
    A multi-artifact detection framework: This work introduces the first multi-artifact detection framework that does not reply on specific video quality assessment models. All the artifact prediction heads in this model are based on the same artifact-aware features, enabling ten artifacts to be detected in a single forward pass of the model, thus significantly improving model efficiency.
  2. 2.
@@
  4. 4.
    Recurrent Memory Vision Transformer (RMViT) module: This is the first time that the recurrent memory mechanism [4] has been used in the context of artifact detection. It has previously been proved to be effective for long sequence processing when employed in language modeling [4].
@@ 2 Related Work
Video quality assessment. Although subjective quality assessment offers a gold standard for measuring the perceived quality of streamed video content, it is impractical for online delivery, where objective quality models are instead used to predict perceptual video quality in a more efficient manner. Early objective video quality assessment methods [45, 47, 33, 46, 44, 36, 53, 30, 28, 29, 35] typically rely on hand-crafted models that are based on classic signal processing theories to exploit different properties of the human visual system. Some of these quality metrics have further been ‘fused’ with other video features in a regression framework [23, 25, 42] to achieve better prediction performance. More recently, many quality models [17, 51, 49, 2, 21, 19] have exploited deep network architectures which can learn from ground-truth subjective data. These models have been further enhanced based on more advanced training strategies [14, 11, 26], which enables the use of more diverse training material without the need to perform expensive subjective tests.
@@ 3 Method
In this framework, each frame of the input video signal is first processed by the Artifact-aware Dynamic Feature Extraction (ADFE) module, which outputs a frame embedding, $\mathbf{h}\in\mathbb{R}^{2048\times 1}$. All the extracted frame embeddings are then fed into the Recurrent Memory Vision Transformer (RMViT) module to obtain a sequence-level representation $\mathbf{v}\in\mathbb{R}^{128\times 1}$. $\mathbf{v}$ is then shared by ten Prediction Heads (with the same network architecture but different model parameters) as their input, each of which outputs a binary label $\varphi_{j}\in\{1,0\},j=1,2\dots\text{ or }10$ to indicate the presence of an artifact type. The network structures, the training database and the model optimization strategy are described in detail below.
@@ 3.2 Training Database
To further randomize the order of artifact introduction, based on the 1200 source patches created, we produced more training patches by generating ten types of artifacts in a random sequence. Each artifact, similar to the baseline database, is associated with a 50% probability of inclusion. This process has been repeated four times, producing another 4,800 training patches with artifacts synthesized in a random order.
@@ 4 Experiment Configuration
5 Results and Discussion
5.1 Overall performance
Table 1 and Table 2 summarize the detection performance of the proposed method compared to the other seven artifact detectors on two test databases. It can be observed that MVAD outperforms all the other methods in each artifact category across the two databases. The detection performance is particularly superior (above 90%) for blockiness, spatial blur and aliasing artifacts, much higher than that of the second best performers. However it is also clear, despite the evident performance improvements, that the proposed method can be further enhanced for challenging artifact cases such as motion blur, graininess and banding. We have plotted the ROC figures for each artifact category in two databases and provided them in Supplementary, which also confirms the effectiveness of the proposed method from a different perspective.
@@ 5.2 Ablation Study
In order to verify the effectiveness of the primary contributions described in section 1, we performed ablation studies to generate four variants of MVAD and compared their performance with the original model. Only the BVI-Artifact database was used in this study (as it contains all ten artifacts types tested). Table 3 provides the results of the ablation study.
Artifact-aware Dynamic Feature Extraction (ADFE). To confirm the contribution of the proposed ADFE module, we replaced it with the feature extractor used in [11, 14] (a pyramid network without the aritifact-aware masking) and obtained (v1). It can be observed that the performance of (V1) is worse than the original MVAD for most artifact classes, which verifies the importance of the proposed ADFE module. In the Supplementary, we have provided a visualization example of the guided mask to demonstrate its influence.
@@ 5.3 Complexity Analysis
The complexity figures of the proposed method and seven benchmark approaches are provided in Table 4. It is noted that MVAD, MaxVQA and VIDMAP can detect multiple visual artifacts; CAMBI, BBAND, EFENet and MLDBD and Wolf et al. are single artifact detectors. We can observe MVAD is associated with a relatively large model size and slow runtime value compared to MaxVQA. This is a limitation of this work and should be addressed in future work.
@@ 6 Conclusion
In this paper, we proposed a novel Multiple Visual Artifact Detector (MVAD) for video streaming. It comprises a new Artifact-aware Dynamic Feature Extractor (ADFE) and a Recurrent Memory Vision Transformer (RMViT) module to capture spatial and temporal information for multiple Prediction Heads. It outputs binary artifact predictions for the presence of ten common visual artifacts. We also developed a large and diverse training dataset based on Adversarial Data Augmentation which has been employed to optimize the proposed model. This multi-artifact detector, MVAD, is the first of its type that does not rely on video quality assessment models. We have demonstrated its superior performance compared to seven existing artifact detection methods when tested on two large benchmark databases.
@@
Future work should focus on model complexity reduction, enhancing MVAD performance on source artifacts such as graininess, banding and motion blur, and improving model generalization for an increased number of diverse artifact types.
Appendix A Appendix
A.1 ROC Plots
Figure 4 and Figure 6 show the ROC plots for eight artifact detection methods on each artifact tested based on the Maxwell and BVI-Artifact database. The ROC (Receiver Operating Characteristic) curves provide a visual representation of the performance of these methods, illustrating the trade-off between the true positive rate (TPR) and the false positive rate (FPR) across different threshold settings.
Figure 4: The ROC curves for different artifact categories in the Maxwell [50] database.
A.2 Visualization of the Guided Mask in ADFE
Figure 5 showcases the influence of the guided mask $\mathbf{M}$ in the ADFE module, which does emphasize the regions with visual artifacts.
Figure 5: Visualization of the guided mask $\mathbf{M}$ generated by the ADFE module.
A.3 Artifacts Synthesis Methods
Table 5 provides a summary of the synthesis methods used to generate various video artifacts. In the Augmented Database, different parameters were employed to produce artifacts with four visibility levels.
Figure 6: The ROC curves for different artifact categories in the BVI-Artifact [10] database.
Table 5: Artifacts synthesis methods used for multiple artifacts generation.
Artifacts Synthesis Methods Levels Parameters
Aliasing Spatial re-sampling [12] Very noticeable Sampling ratio=4
Noticeable Sampling ratio=3
Subtle Sampling ratio=2
Very subtle Sampling ratio=1.5
Banding Quantization [12] Very noticeable Quantization ratio=5
Noticeable Quantization ratio=4
Subtle Quantization ratio=3
Very subtle Quantization ratio=2
Dark Scene Brightness/contrast adjustment [13] Very noticeable Decrease ratio=4
Noticeable Decrease ratio=3
Subtle Decrease ratio=2
Very subtle Decrease ratio=1.5
Motion Blur Convolution (motion blur) [37] Very noticeable Frame numbers=16
Noticeable Frame numbers=12
Subtle Frame numbers=8
Very subtle Frame numbers=4
Graininess Gaussian noise [12] Very noticeable standard deviation=50
Noticeable standard deviation=25
Subtle standard deviation=10
Very subtle standard deviation=5
Blockiness JPEG compression [12] Very noticeable Quantization parameter= 47
Noticeable Quantization parameter= 42
Subtle Quantization parameter= 37
Very subtle Quantization parameter= 32
Frame Drop Random frame dropping [48] Very noticeable Frame length = 16
Noticeable Frame length = 12
Subtle Frame length = 8
Very subtle Frame length = 4
Spatial Blur Convolution (Gaussian blur) [12] Very noticeable Kernel size=9
Noticeable Kernel size=7
Subtle Kernel size=5
Very subtle Kernel size=3
Transmission Error Packet loss and error concealment [12] Very noticeable bitstream filter = 4M
Noticeable bitstream filter = 2M
Subtle bitstream filter = 1M
Very subtle bitstream filter = 0.5M
Black Screen Random black frames replacement Very noticeable Frame length = 16
Noticeable Frame length = 12
Subtle Frame length = 8
Very subtle Frame length = 4
A.4 Broader Impacts
The development and implementation of the MVAD framework for detecting multiple visual artifacts in streamed videos have the following impacts.
On the positive side, the MVAD system supports automatic inspection of video content to identify visual artifacts, significantly reducing labor costs and increasing the efficiency of quality control processes in video streaming services. MVAD can also potential improve the quality of streamed videos by incorporating with video enhancement methods, leading to a better viewing experience for end users.
The implementation of MVAD can also result in negative impacts. The training and deployment of sophisticated machine learning models such as MVAD require substantial computational resources, which can lead to increased energy consumption and negative environmental impact. Further model complexity reduction can alleviate this issue, and it remains our future work.
A.5 License of Code and Data
Table 6 summarizes the license associated with the code and data used and generated in this work.
Table 6: License information for the Code and datasets used and generated in this work.
Code/Data Size Dataset URL License/Terms of Use
Training Datasets
BVI-DVC [24] 800 https://fan-aaron-zhang.github.io/BVI-DVC/ Academic research.
BVI-CC [15] 90 https://fan-aaron-zhang.github.io/BVI-CC/ Academic research.
NFLX-public [23] 70 https://github.com/Netflix/vmaf/blob/master/resource/doc/datasets.md Academic research.
LIVE-HFR [27] 88 https://fan-aaron-zhang.github.io/BVI-HFR/ Academic research.
Adobe240 [39] 133 https://www.cs.ubc.ca/labs/imager/tr/2017/DeepVideoDeblurring/ MIT license.
Test Datasets
BVI-Artifact [10] 480 https://chenfeng-bristol.github.io/BVI-Artefact/ Academic research.
Maxwell [50] 4,543 https://github.com/VQAssessment/ExplainableVQA MIT license.
Code of Benchmark Methods
MaxVQA [50] - https://github.com/VQAssessment/ExplainableVQA MIT license.
VIDMAP [12] - https://github.com/utlive/VIDMAP Academic research.
CAMBI [40] - https://github.com/Netflix/vmaf/blob/master/resource/doc/cambi.md BSD+Patent.
BBAND [41] - https://github.com/google/bband-adaband Apache-2.0 license.
EFENet [57] - https://github.com/wdzhao123/DENets No licence.
MLDBD [58] - https://github.com/wdzhao123/MLDBD No licence.
Wolf el at. [48] - https://its.ntia.gov/umbraco/surface/download/publication?reportNumber=TR-09-456.pdf Academic research.
Code and Data generated in this work
MVAD - https://github.com CC-BY-4.0
Training data [24] 50,800 https://github.com Academic research.

### v2

@@ Abstract
Visual artifacts are often introduced into streamed video content, due to prevailing conditions during content production and delivery. Since these can degrade the quality of the user’s experience, it is important to automatically and accurately detect them in order to enable effective quality measurement and enhancement. Existing detection methods often focus on a single type of artifact and/or determine the presence of an artifact through thresholding objective quality indices. Such approaches have been reported to offer inconsistent prediction performance and are also impractical for real-world applications where multiple artifacts co-exist and interact. In this paper, we propose a Multiple Visual Artifact Detector, MVAD, for video streaming which, for the first time, is able to detect multiple artifacts using a single framework that is not reliant on video quality assessment models. Our approach employs a new Artifact-aware Dynamic Feature Extractor (ADFE) to obtain artifact-relevant spatial features within each frame for multiple artifact types. The extracted features are further processed by a Recurrent Memory Vision Transformer (RMViT) module, which captures both short-term and long-term temporal information within the input video. The proposed network architecture is optimized in an end-to-end manner based on a new, large and diverse training database that is generated by simulating the video streaming pipeline and based on Adversarial Data Augmentation. This model has been evaluated on two video artifact databases, Maxwell and BVI-Artifact, and achieves consistent and improved prediction results for ten target visual artifacts when compared to seven existing single and multiple artifact detectors. The source code and training database will be available at https://chenfeng-bristol.github.io/MVAD/.
@@ 1 Introduction
With the significant growth in subscribers to streaming services [39], such as Netflix and Amazon Prime Video, video streaming applications are now one of the largest consumers of global Internet bandwidth. For video streaming service providers, it is essential to monitor the quality of a user’s experience during viewing. However, this process is complex because streamed content can be impacted by multiple visual artifacts introduced at different production and delivery stages including acquisition, post-production, compression, and transmission [7]. For example, source artifacts such as motion blur, darkness and graininess can be produced during acquisition; banding and aliasing may be introduced in post-production; video compression can introduce spatial blur and blockiness; and finally packet loss in transmission leads to transmission errors, frame dropping, or black frames. These artifacts can affect the perceived quality of streamed video content and thus degrade the quality of the user’s experience. Therefore, it is important to identify the presence of these artifacts in order to enable appropriate enhancement processes and provide system feedback.
Although video quality assessment is a well-established research area, with numerous classic [47, 49, 55, 31, 29, 30, 36] and learning-based approaches [24, 20, 22, 43, 12] developed over the past two decades, these methods can only provide a quantitative prediction of perceptual quality, rather than detect and identify the artifacts appearing in a video sequence. To achieve accurate artifact detection, some existing works apply thresholding on predicted quality scores to determine the existence of artifacts, with MaxVQA [52] being a typical example. Other works have developed bespoke models for individual artifacts, such as CAMBI [41], BBAND [42], EFENet [59], MLDBD [60] and [50]. This latter approach is a less practical solution for real-world scenarios where multiple artifacts co-exist and interact. In a recent benchmarking experiment [11], all these methods were reported to perform poorly on a video database containing multiple source and non-source artifacts - with the average artifact detection accuracy for the best performers being only 55%.
@@
  • 1)
    A multi-artifact detection framework: This work introduces the first multi-artifact detection framework that does not rely on specific video quality assessment models. All the artifact prediction heads in this model are based on the same artifact-aware features, enabling ten artifacts to be detected in a single forward pass of the model, thus significantly improving model efficiency.
  • 2)
@@
  • 4)
    Recurrent Memory Vision Transformer (RMViT) module: The RMViT module demonstrates superior performance in capturing the temporal characteristics of artifacts compared to other pooling methods. This is the first time that the recurrent memory mechanism [5] has been used in the context of artifact detection. It has previously been proved to be effective for long sequence processing when employed in language modeling [5].
@@ 2 Related Work
Video quality assessment. Although subjective quality assessment offers a gold standard for measuring the perceived quality of streamed video content, it is impractical for online delivery, where objective quality models are instead used to predict perceptual video quality in a more efficient manner. Early objective video quality assessment methods [47, 49, 34, 48, 45, 37, 55, 31, 29, 30, 36] typically rely on hand-crafted models that are based on classic signal processing theories to exploit different properties of the human visual system. Some of these quality metrics have further been ‘fused’ with other video features in a regression framework [24, 26, 43] to achieve better prediction performance. More recently, many quality models [18, 53, 51, 2, 22, 20, 46] have exploited deep network architectures which can learn from subjective data. These models have been further enhanced based on more advanced training strategies [15, 12, 27], which enables the use of more diverse training material without the need to perform expensive subjective tests.
@@ 3 Method
In this framework, each frame of the input video signal is first processed by the Artifact-aware Dynamic Feature Extraction (ADFE) module, which outputs a frame embedding, $\mathbf{h}\in\mathbb{R}^{2048\times 1}$. All extracted frame embeddings are fed into the Recurrent Memory Vision Transformer (RMViT) module to obtain a sequence-level representation $\mathbf{v}\in\mathbb{R}^{128\times 1}$. $\mathbf{v}$ is then shared by ten Prediction Heads (with the same network architecture but different model parameters) as input, each of which outputs a binary label $\varphi_{j}\in\{1,0\},j=1,2\dots\text{ or }10$ to indicate the presence of an artifact type. The network structures, the training database and the model optimization strategy are described in detail below.
@@ 3.2 Training Database
To further randomize the order of artifact introduction, based on the 1200 source patches created, we produced more training patches by generating ten types of artifacts in a random sequence. Each artifact, similar to the baseline database, is associated with a 50% probability of inclusion. This process was repeated four times, producing 4,800 additional training patches with artifacts synthesized in random order.
@@ 4 Experiment Configuration
5 Results and Discussion
5.1 Overall performance
Table 1 and Table 2 summarize the detection performance of the proposed method compared to the other seven artifact detectors on two test databases. It can be observed that MVAD outperforms all the other methods in each artifact category across the two databases. The detection performance is particularly superior (above 90%) for blockiness, spatial blur and aliasing artifacts, much higher than that of the second best performers. However it is also clear, despite the evident performance improvements, that the proposed method can be further enhanced for challenging artifact cases such as motion blur, graininess and banding. We have plotted the ROC figures for each artifact category in two databases and provided in Supplementary, which also confirms the effectiveness of the proposed method from a different perspective.
@@ 5.2 Ablation Study
In order to verify the effectiveness of the primary contributions described in section 1, we performed ablation studies to generate four variants of MVAD and compared their performance with the original model, results shown in Table 3. Only the BVI-Artifact database was used in this study (as it contains all ten artifact types tested).
Artifact-aware Dynamic Feature Extraction (ADFE). To confirm the contribution of the proposed ADFE module, we replaced it with the feature extractor used in [12, 15] (a pyramid network without the aritifact-aware masking) and obtained (v1). It can be observed that the performance of (v1) is worse than the original MVAD for most artifact classes, which verifies the importance of the proposed ADFE module. In the Supplementary, we have provided a visualization example of the guided mask to demonstrate its influence.
@@ 5.3 Complexity Analysis
The complexity figures of the proposed method and seven benchmark approaches are provided in Table 4. It is noted that MVAD, MaxVQA and VIDMAP can detect multiple visual artifacts; CAMBI, BBAND, EFENet and MLDBD and Wolf et al. are single artifact detectors. Among these models, MVAD has a relatively large model size and slow runtime value compared to MaxVQA.
@@ 6 Conclusion
In this paper, we proposed a novel Multiple Visual Artifact Detector (MVAD) for video streaming. It comprises a new Artifact-aware Dynamic Feature Extractor (ADFE) and a Recurrent Memory Vision Transformer (RMViT) module to capture spatial and temporal information for multiple Prediction Heads. It outputs binary artifact predictions for the presence of ten common visual artifacts. We also developed a large and diverse training dataset based on Adversarial Data Augmentation to optimize the proposed model. This multi-artifact detector, MVAD, is the first of its type that does not rely on video quality assessment models. We demonstrated its superior performance compared to seven existing artifact detection methods on two large benchmark databases.
@@
Future work should focus on reducing model complexity using knowledge distillation [14] and model compression [4] technologies, enhancing MVAD performance on source artifacts such as graininess, banding and motion blur, and improving the generalization of the model for an increasing number of different artifact types.
Acknowledgement. The authors appreciate the funding from Amazon Research Award, Fall 2022 CFP and the UKRI MyWorld Strength in Places Programme (SIPF00006/1), the University of Bristol. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of Amazon.